# Decoding Visual Experience and Mapping Semantics through Whole-Brain Analysis Using fMRI Foundation Models

### v2

@@ A.2 Method
For our training process, we consistently utilized a seed value of 42. We employed an Adam optimizer, setting the learning rate at $1\times 10^{-4}$. The decay rates were configured differently for various components of the training: $0.03$ for the contrastive learning part and $0.02$ for the diffusion model training. The number of training steps was set to 50,000 for the contrastive learning phase and 200,000 for the diffusion model phase, for each individual model. In the case of universal model training, which involved data from four subjects, we quadrupled the training steps to accommodate the increased data volume.
@@
The self-supervised clustering method that we used is k-clustering after the encoded image features by CLIP. It is not easy to determine how many clusters we should use in the analysis. The datasets include broad range of image types and classes, and we also conducted the elbow analysis for better cluster numbers. As we showed in Fig. 7, the sum of the squared distance is decreasing with larger cluster number. But there is no obvious elbow curve occurs even in 20 clusters. For better clustering analysis purpose, we used 5 clusters for our study.
@@
For the evaluation and analysis of our model, we employed various metrics, including the mean cosine distance between scenario text and training images. This metric is computed as described in Equation 3. We provide the Python scipy code [59] used for calculating the mean cosine distance for each scenario texts, as follows:
@@
cosine_distances = cdist(image_embeddings, _gt_text_embeddings, metric=’cosine’)

### v3

@@ A.2 Method
fMRI Foundation Model.
To leverage large-scale neural representations, we utilize the fMRI foundation model developed by Thomas et al. Thomas et al. [2022], which employs a decoder-only Transformer architecture (GPT-2) Radford et al. [2019]. To standardize inputs and capture whole-brain responses efficiently, fMRI data is parcellated into 1024 regions of interest using the DiFuMo1024 atlas Dadi et al. [2020]. This fine-grained soft parcellation preserves approximately 70% of the explained variance in neural signals while reducing spatial dimensionality. Formally, an fMRI sequence is defined as $X\in\mathbb{R}^{T\times N}$, where $T$ represents time points and $N=1024$ represents the network parcels. During pre-training, the model treats each fMRI volume (TR) as a discrete token. Given different fMRI datasets have varies TRs, the fMRI foundation model treats TR as positional embeddings added to the fMRI tokens. These inputs are projected into an embedding space $E\in\mathbb{R}^{T\times D}$, where $D$ is the embedding dimension. The model is trained via a self-supervised autoregressive objective to predict the fMRI image at $T+1$ based on the history of previous fMRI images, effectively learning generalizable temporal brain dynamics. You can find details such as acquisition protocols in original paper. We initialize our encoder with these pre-trained weights [^3] Code available at: https://github.com/athms/learning-from-brains and fine-tune the model on the BOLD5000 dataset for the visual decoding task.
@@
For our training process, we consistently utilized a seed value of 42. We employed an Adam optimizer, setting the learning rate at $1\times 10^{-4}$. The decay rates were configured differently for various components of the training: $0.03$ for the contrastive learning part and $0.02$ for the diffusion model training. The number of training steps was set to 50,000 for the contrastive learning phase and 200,000 for the diffusion model phase, for each individual model. In the case of universal model training, which involved data from four subjects, we quadrupled the training steps to accommodate the increased data volume. We also conducted hyperparameters search as described in Table 10.
Table 10: WAVE Model Training Hyperparameters. Configuration details for training phases across individual and universal model settings.
Hyperparameter Value
Random Seed 42
Optimizer Adam
Learning Rate Log-Uniform(5e-5, 1e-3)
Weight Decay
    Contrastive Phase Log-Uniform(1e-3, 5e-2)
    Diffusion Phase Log-Uniform(1e-3, 5e-2)
@@
The self-supervised clustering method that we used is k-clustering after the encoded image features by CLIP. It is not easy to determine how many clusters we should use in the analysis. BOLD5000 includes broad range of image types and classes, and we also conducted the elbow analysis for better cluster numbers. As we showed in Fig. 7, the sum of the squared distance is decreasing with larger cluster number. But there is no obvious elbow curve occurs even in 20 clusters. For better clustering analysis purpose, we used 5 clusters for our study.
@@
For the evaluation and analysis of our model, we employed various metrics, including the mean cosine distance between scenario text and training images. This metric is computed as described in Equation 3. We provide the Python scipy code [Virtanen et al., 2020] used for calculating the mean cosine distance for each scenario texts, as follows:
@@
cosine_distances = cdist(image_embeddings,
                    _gt_text_embeddings, metric=’cosine’)